OpenStudy (anonymous):

System of equation: substitution method 8x+3y=26 2x=y-4

OpenStudy (anonymous):

\[x = \frac{ y }{ 2 } - 2\] \[8(\frac{ y }{ 2 } - 2) + 3y = 26\] \[4y - 16 + 3y = 26\] \[7y - 16 = 26\] \[7y = 44\] \[y = 6.29\] Are you supposed to have a decimal? If not, y = 10 hehe

OpenStudy (anonymous):

no your trying to find x and y for cordinates

OpenStudy (anonymous):

x = 7

OpenStudy (anonymous):

how did u get 7 all i see is the y

OpenStudy (texaschic101):

or... 2x = y - 4 2x + 4 = y now sub 2x + 4 in for y in the 1st equation 8x + 3y = 26 8x + 3(2x + 4) = 26 8x + 6x + 12 = 26 14x = 26 - 12 14x = 14 x = 1 now sub 1 in for x in either of the original equations 2x = y - 4 2(1) = y - 4 2 = y - 4 2 + 4 = y 6 = y now lets check... 8x + 3y = 26 8(1) + 3(6) = 26 8 + 18 = 26 26 = 26 (correct) x = 1 and y = 6

OpenStudy (anonymous):

Well, using a different method, I got x = 7, y = 10. I'm not that good with this method, can anyone else correct me.

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

...I keep doing the wrong variable first...

OpenStudy (texaschic101):

it doesn't matter which variable you do first

OpenStudy (anonymous):

Well, can you solve for x first?

OpenStudy (texaschic101):

I did y first because it was already by itself

OpenStudy (texaschic101):

2x = y - 4 x = (y - 4)/2 x = 1/2y - 2 now sub 1/2y - 2 in for x 8x + 3y = 26 8(1/2y - 2) + 3y = 26 4y - 16 + 3y = 26 7y = 26 + 16 7y = 42 y = 6 now sub 6 in for y 2x = y - 4 2x = 6 - 4 2x = 2 x = 1

OpenStudy (texaschic101):

it does not matter which variable you solve for first, as you can see, you will get the same answer

OpenStudy (anonymous):

i just wanna check if i got it right x-7y=13 3x-5y=23 for this equation i got {6,-1}

OpenStudy (texaschic101):

x - 7y = 13 --> x = 7y + 13 3x - 5y = 23 3(7y + 13) - 5y = 23 21y + 39 - 5y = 23 16y = -16 y = -1 x - 7y = 13 x -7(-1) = 13 x + 7 = 13 x = 13 - 7 x = 6 you are correct :) if your ever in doubt, you can check you answers by subbing them back into the equations. 3x - 5y = 23 3(6) - 5(-1) = 23 18 + 5 = 23 23 = 23 (correct) if they come out equal, then your answers are correct....if they don't come out equal, then there is a problem.

OpenStudy (anonymous):

can you use any of both equations to double check

OpenStudy (texaschic101):

you can use either one or both

OpenStudy (anonymous):

okay thank you :)

OpenStudy (texaschic101):

your very welcome :)